PLEASE Help!!!
Factoring to solve Quadratic Equations.
Use the Zero-Product Property to solve.
1. (x + 7) (4x - 5) = 0
2. (2x + 7) (2x - 7) = 0
3. (5x + 6) (4x + 5) = 0

Answer by checkley71(8403)

(X+7)(X-5)=0
X+7=0
X=-7 ANSWER.
X-5=0
X=5 ANSWER.
---------------------------
2X+7)(2X-7)=0
2X+7=0
2X=-7
X=-7/2 ANSWER.
2X-7=0
2X=7
X=7/2 ANSWER.
---------------------------------
(5X+6)(4X+5)=0
5X+6=0
5X=-6
X=-6/5 ANSWER.
4X+5=0
4X=-5
X=-5/4 ANSWER.